I need help in evaluating the following expressions: Thanks in advance.
1. 5.4 - (-6.9)
2. (13 - 5) - (-2 + 6)
I also need to simplify the following:
1. -19 - x - 22
2. 30x*(1/6)

Answer by checkley71(8403)

5.4-(-6.9)=5.4+6.9=12.3
-------------------------------------
13-5-(-2+6)=8+2-6=4
------------------------------------
-19-x-22
-41-x
----------------------------------------------
30x*1/6
5x
------------------------------------------